OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0). y − 4 = −one half(x + 6) y − 4 = 2(x + 6) y + 6 = −one half(x − 4) y + 6 = 2(x − 4) @Isaiah.Feynman

OpenStudy (anonymous):

@canadalover_

OpenStudy (bakonloverk):

Do you know how to do it?

OpenStudy (anonymous):

the hell if i now this sht (I have been doning this for 2 hours now)

OpenStudy (bakonloverk):

y-y1=m(x-x1) all you have to do is input the numbers

OpenStudy (bakonloverk):

find the slope first

OpenStudy (anonymous):

ok

OpenStudy (bakonloverk):

do you know how to do that?

OpenStudy (anonymous):

y - 4 = -one half (x + 6) ......................Answer y - 4 = 2(x + 6) y + 6 = -one half (x - 4) y + 6 = 2(x - 4) @Kstate